Given that P (2,-3) lies on the line kx+3y+9=0, find K
Answer by Alan3354(69443)

Given that P (2,-3) lies on the line kx+3y+9=0, find K
------------
kx+3y+9 = 0
kx = -3y - 9 = -3*-3 - 9 = 0
2k = 0
k = 0